A galvanometer of resistance 100 ohm is converted to a voltmeter of range 10 V by connecting a resistance of 10 kohm .The resistance required to convert the same galvanometer to an ammeter of range 1A is

Grade:12

1 Answers

Khimraj
5 years ago
RG= 100 ohm
IG = 10/10000 = 1mA
I = 1A
(I – IG)Rs = IG*RG
(1 – 0.001)Rs = 10-3*100
SO Rs = 0.1 ohm.
Hope it clears.